OpenStudy (anonymous):

find the smallest and largest values of the function f(r) = r - sin2r on the interval [0,π]

OpenStudy (anonymous):

this one is easy

OpenStudy (anonymous):

f'(r)= 1- 2 cos (2r)

OpenStudy (anonymous):

and it equals to 0

OpenStudy (anonymous):

1-2 cos(2r)=0 2 cos(2r)=1 cos(2r)=1/2 cos (pi/3) =1/2 thus r= pi/6

OpenStudy (anonymous):

check left and right

OpenStudy (anonymous):

hence cos2r = 1/2 so that \[r = \pi / 6 or r = \pi - (\pi/6) or 5\pi/6\]

OpenStudy (anonymous):

mm but i guess it can be pi - (pi/6) and 5pi/6 too

OpenStudy (anonymous):

f'(r)= 1- 2 cos (2r) less than pi/6 , it is negative more than pi/6 , it is positive

OpenStudy (anonymous):

the fact that slope change from negative to positive mean pi/3 is min

OpenStudy (anonymous):

hmm k

OpenStudy (anonymous):

oops, I meant pi/6

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

A plot is attached.

OpenStudy (anonymous):

thanks :D